OpenStudy (anonymous):

what is the integral of the function 1/x^7 from 1 to infinity? I already know it converges to a number because of p = 7 is greater than 1

OpenStudy (anonymous):

hello!

OpenStudy (anonymous):

well.. first find the integral .. which is (-1/6)(1/x^6).. then apply limits..

OpenStudy (amistre64):

improper integrals eh ..... havent tried them out yet ..

OpenStudy (anonymous):

well, with the integral of the function, it looks like what ever n could be from 1 to infinity, it would be too small to change the value of (-1/6), so maybe the answer is -1/6. I will try it

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

answer is +(1/6)

OpenStudy (anonymous):

you would get.. 0 - (-1/6) = 1/6

OpenStudy (anonymous):

oh, whoops, math error! thanks

OpenStudy (amistre64):

x^-7 ints up to x^-6/-6 \[\frac{-1}{6(inf)^6}-\frac{-1}{6(1)^6}=0-(-1/6)=1/6\]

OpenStudy (anonymous):

okay, now i see it better

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (zarkon):

These integral should be written \[\int\limits_{1}^{\infty}\frac{1}{x^7}dx=\lim_{t\to\infty}\int\limits_{1}^{t}\frac{1}{x^7}dx\] Integrate like you normally would using the FTC then take the limit